# A Case Study on Energy-Efficient Edge AI Crack Segmentation

### 15 Apr 2026 v1

@@ IV-B Custom Hardware Implementation on FPGA
The FPGA implementation is built upon a High-level Synthesis (HLS) hardware library that provides custom hardware architectures for a wide range of Deep Neural Network (DNN) layers and components. These architectures are highly customizable, enabling the implementation of diverse neural topologies and facilitating rapid design-space exploration. The hardware library is implemented as a collection of C++ template functions equipped with HLS annotations. The overall architecture is optimized for low-power consumption and ultra-low latency operation. This is primarily accomplished by storing all weights and the majority of intermediate results in on-chip memory, as off-chip data transfers incur higher energy costs and introduce additional latency. Nevertheless, a design relying solely on on-chip resources inherently limits the size of DNN models that can be feasibly implemented. To overcome this constraint, we developed a hybrid hardware architecture that supports off-chip implementation of skip and residual connections, effectively relaxing the limitations on model size and structural complexity. The architecture is fully pipelined, allowing all layers to operate concurrently and initiate computation immediately once their inputs become available, thereby reducing both latency and energy consumption. Individual hardware modules corresponding to each layer are interconnected via on-chip data streams and assembled into a single top-level module, as illustrated in Fig. 1.
@@ V Results
V-A Accuracy–Efficiency Trade-off Analysis
@@
Considering the model results without knowledge distillation, the original U-Net ($c=32$) serves as the baseline. Reducing the base channel count to $c=16$ reduces the number of Multiply-Accumulate (MAC) operations by approximately $4\times$, from 24.21G to 6.10G, while the accuracy drop is negligible - only 0.18pps in weighted IoU and 0.08pps in mean IoU. Reducing further to $c=8$ provides an additional $\sim 4\times$ MAC reduction to approximately 1.55G, with a still modest accuracy cost of only 3.63pps in weighted IoU and 2.11pps in mean IoU relative to the baseline - a competitive result given the overall $16\times$ reduction in computational cost. Scaling down to $c=2$, however, incurs steep accuracy costs: a drop of 13.60pps in weighted IoU and 10.46pps in mean IoU compared to $c=16$, confirming that $c=16$ and $c=8$ represent the best efficiency–accuracy trade-offs in this family. As expected, SegFormer B5 outperforms all UNet variants without knowledge distillation, owing to its considerably larger capacity and modern transformer-based architecture.
With knowledge distillation, all UNet variants improve. The baseline model ($c=32$) achieves a weighted IoU of $51.06\%$ and a mean IoU of $71.92\%$—within 2.07pps of the teacher SegFormer B5 ($53.13\%$ weighted IoU), showing that KD brings the student nearly to the teacher’s level. Crucially, the efficiency gains identified without KD are preserved: $c=16$ achieves $50.63\%$ weighted IoU and $71.74\%$ mean IoU, only 0.43 and 0.18pps below $c=32$, while still benefiting from the same $4\times$ MAC reduction. Similarly, $c=8$ achieves $49.25\%$ weighted IoU and $70.96\%$ mean IoU—only 1.81pps and 0.96pps below the best UNet model —at a $16\times$ overall reduction in compute.
Overall, these results show that KD not only improves segmentation results across all UNet variants, but also makes the models significantly more robust to channel reduction: the mean IoU gap between $c=32$ and $c=8$ shrinks from 2.19pps without KD to only 0.96pps with KD, which shows that KD compensates for the capacity loss introduced by pruning. In the next subsection, we evaluate each variant across several edge platforms under quantization.
TABLE I: Comparison of U-Net variants with different channel scaling factors, with and without knowledge distillation (KD). The Hybrid-Segmentor model was proposed by Goo et al [11]. SegFormer B5 is included as a high-capacity reference.
@@
V-B Hardware Implementation Results
Table IV shows the results for knowledge-distilled models across three embedded platforms and the FPGA, reporting segmentation accuracy (weighted and mean IoU), throughput (FPS), and power and energy metrics. In addition to idle and runtime power, both dynamic and runtime energy efficiency (frames/J) are provided. Our analysis focuses primarily on dynamic energy efficiency, as it isolates the energy consumption attributable to active inference. In contrast, runtime energy efficiency is affected by background processes and system-level overhead, making cross-platform comparisons less reliable. Fig. 2 further illustrates the trade-off between dynamic energy efficiency and mean IoU across the evaluated configurations.
V-B1 FPGA results
For an efficient FPGA implementation PTQ is used to reduce the precision of the model parameters. Two experiments were conducted; in the first experiment, the weights and activations were quantized to int8 as shown in Table II this experiment achieved a similar or higher IoU compared to the fp32 experiment due to regularization. In the second experiment, the weights were quantized to int4, which resulted in a considerable drop in accuracy as shown in Table II. We attribute this drop to PTQ and expect quantization aware training (QAT) to recover this drop. Based on the results in Table II int8 models were chosen for implementation on FPGA, for which the AMD Vitis 2024.2 HLS tool was used and the target platform was set to TySOM-2A-ZU19EG. Table III presents the resources required to implement the chosen models on the target platform. Two experiments were conducted; in the first experiment, the resource-heavy skip connections of the UNet model were kept on chip, while in the second experiment, they were moved to off-chip memory. This made it possible to implement the larger model with $c=16$. Skip connections were moved to off-chip memory by creating an AXI interface for each skip connection, which is used for reading and writing. This is required due to the high read and write speeds and the fact that all skip connections are accessed in parallel. On-chip buffers were added to the design to bridge the gap between the high-speed PL and the slow off-chip memory. The extra buffers and interfaces lead to an average increase of $2.19\%$ and $1.60\%$ in LookUp Tables (LUTs) and Flip-Flops (FFs) respectively, which is within the acceptable overhead. This shows the tradeoff between moving skip connections off chip and keeping them on chip, which is less efficient for small models i.e., with $c=2$. Table IV shows the throughput, IoU, energy and efficiency of the implemented models on the FPGA. All FPGA implementations achieved the same IoU as the corresponding software implementations.
V-B2 Edge Computing Platforms
@@
Figure 2: Pareto frontier showing the trade-off between accuracy (weighted IoU) and energy efficiency across different platforms and model sizes. FPGA on-chip: models UNet skip connections stored on chip, FPGA off-chip: models UNet skip connections stored off chip in the memory
@@ VI Conclusion and Future Work
In this work, we studied crack segmentation on resource-constrained edge platforms from two sides: The model design and the target hardware. For this, we scaled the U-Net architecture over three orders of magnitude in parameter count and analyzed the trade-off between IoU and computational cost. Our results show that reducing the model size can strongly lower the computational cost, while some of the smaller variants still remain close to the baseline in terms of segmentation accuracy. To improve our scaled UNet models further, we applied knowledge distillation by using SegFormer B5 as the teacher model. This improved all U-Net variants.
We then evaluated all variants on several edge platforms and compared their throughput, power consumption, and energy efficiency. We also developed a custom FPGA-based architecture for single-batch inference. Among all tested platforms, the FPGA (on-chip) implementation with the UNet Base 4 using KD and int8 quantization achieves an energy efficiency of 204.99 Frames/J and a throughput of 398 FPS, while improving the mean IoU by 2.79pps compared to the baseline model (UNet Base 32 without KD). Compared to the Hybrid-Segmentor model from Goo et al. [11], which is - as far as we know - currently the best tested model on this dataset, our aforementioned UNet Base 4 model with KD is 6.32pps better in mean IoU.

### 06 May 2026 v2

@@ IV-B Custom Hardware Implementation on FPGA
Our FPGA implementation uses an High-level Synthesis (HLS) hardware library that provides custom hardware architectures for many common Deep Neural Network (DNN) layers and components. The library is built as a set of C++ template functions with HLS annotations, which makes it easy to support different network topologies and to explore the design space quickly. We design the architecture for low power and low latency. To this end, we keep all weights and most intermediate results in on-chip memory, since off-chip transfers cost more energy and add latency.
Nevertheless, a design relying solely on on-chip resources inherently limits the size of DNN models that can be feasibly implemented. To overcome this constraint, we developed a hybrid hardware architecture that supports off-chip implementation of skip and residual connections, effectively relaxing the limitations on model size and structural complexity. The architecture is fully pipelined, allowing all layers to operate concurrently and initiate computation immediately once their inputs become available, thereby reducing both latency and energy consumption. Individual hardware modules corresponding to each layer are interconnected via on-chip data streams and assembled into a single top-level module, as illustrated in Fig. 1.
@@ V Results
V-A Scope and Fairness of the Cross-Platform Comparison
The comparison in Table IV is a practical evaluation rather than a strictly controlled evaluation. The used models differ across platforms in three ways, each of which is due more to platform-specific constraints rather than to the design itself. First, models targeted at the Coral Edge TPU replace transposed convolutions with nearest-neighbor upsampling to avoid the accuracy drop we observed on this device; All other platforms retain the original transposed-convolution decoder. Second, the numerical precision is not uniform: the RPi5 runs fp32 weights and activations, the Jetson Orin Nano is evaluated in both fp32 and a mixed int8-weights/fp32-activations configuration (native to TensorRT), the Edge TPU and the FPGA executes fully int8. Third, the measurement instrumentation differs between platforms. We therefore interpret the numbers as realistic deployment results for each platform in its best-supported configuration, rather than as an apples-to-apples comparison of the same model across devices.
V-B Accuracy-Efficiency Trade-off Analysis
@@
Considering the model results without knowledge distillation, the original U-Net ($c=32$) serves as the baseline. Reducing the base channel count to $c=16$ reduces the number of Multiply-Accumulate (MAC) operations by approximately $4\times$, from 24.21G to 6.10G, while the accuracy drop is negligible - only 0.18pps in weighted IoU and 0.08pps in mean IoU. Reducing further to $c=8$ provides an additional $\sim 4\times$ MAC reduction to approximately 1.55G, with a still modest accuracy cost of only 3.63pps in weighted IoU and 2.11pps in mean IoU relative to the baseline - a competitive result given the overall $16\times$ reduction in computational cost. Scaling down to $c=2$, however, incurs steep accuracy costs: a drop of 13.60pps in weighted IoU and 10.46pps in mean IoU compared to $c=16$, confirming that $c=16$ and $c=8$ represent the best efficiency–accuracy trade-offs in this family. As expected, SegFormer B5 outperforms all U-Net variants without knowledge distillation, owing to its considerably larger capacity and modern transformer-based architecture.
With knowledge distillation, all U-Net variants improve. The baseline model ($c=32$) achieves a weighted IoU of $51.06\%$ and a mean IoU of $71.92\%$—within 2.07pps of the teacher SegFormer B5 ($53.13\%$ weighted IoU), showing that KD brings the student nearly to the teacher’s level. Crucially, the efficiency gains identified without KD are preserved: $c=16$ achieves $50.63\%$ weighted IoU and $71.74\%$ mean IoU, only 0.43 and 0.18pps below $c=32$, while still benefiting from the same $4\times$ MAC reduction. Similarly, $c=8$ achieves $49.25\%$ weighted IoU and $70.96\%$ mean IoU—only 1.81pps and 0.96pps below the best U-Net model —at a $16\times$ overall reduction in compute.
Overall, these results show that KD not only improves segmentation results across all U-Net variants, but also makes the models significantly more robust to channel reduction: the mean IoU gap between $c=32$ and $c=8$ shrinks from 2.19pps without KD to only 0.96pps with KD, which shows that KD compensates for the capacity loss introduced by pruning. In the next subsection, we evaluate each variant across several edge platforms under quantization.
TABLE I: Comparison of U-Net variants with different channel scaling factors, with and without knowledge distillation (KD). The Hybrid-Segmentor model was proposed in [12]. SegFormer B5 is included as a high-capacity reference.
@@
V-C Hardware Implementation Results
As discussed in Section V-A, the configurations per platform are not identical, the comparison reflects the best-supported deployment path on each device. Table IV shows the results for knowledge-distilled models across three edge platforms and the FPGA, reporting segmentation accuracy (weighted and mean IoU), throughput (FPS), and power and energy metrics. In addition to idle and runtime power, both dynamic and runtime energy efficiency (Frames/J) are provided. Our analysis focuses primarily on dynamic energy efficiency, as it isolates the energy consumption attributable to active inference. In contrast, runtime energy efficiency is affected by background processes and system-level overhead, making cross-platform comparisons less reliable. Runtime energy efficiency does not solve all measurement issues, but it gives a more realistic picture of how the models behave in deployment. Fig. 2 further illustrates the trade-off between dynamic energy efficiency and mean IoU across the evaluated configurations.
V-C1 FPGA results
For an efficient FPGA implementation PTQ is used to reduce the precision of the model parameters. Two experiments were conducted; in the first experiment, the weights and activations were quantized to int8 as shown in Table II this experiment achieved a similar or higher IoU compared to the fp32 experiment due to regularization. In the second experiment, the weights were quantized to int4, which resulted in a considerable drop in accuracy as shown in Table II. We see this drop as a limit of PTQ at very low bit widths and leave a follow-up with quantization aware training (QAT) to future work. Based on the results in Table II int8 models were chosen for implementation on FPGA, for which the AMD Vitis 2024.2 HLS tool was used and the target platform was set to TySOM-2A-ZU19EG. Table III presents the resources required to implement the chosen models on the target platform. Two experiments were conducted; in the first experiment, the resource-heavy skip connections of the U-Net model were kept on chip, while in the second experiment, they were moved to off-chip memory. This made it possible to implement the larger model with $c=16$. Skip connections were moved to off-chip memory by creating an AXI interface for each skip connection, which is used for writing and reading. This is required due to the high write and read speeds and the fact that all skip connections are accessed in parallel. On-chip buffers were added to the design to bridge the gap between the high-speed PL and the slow off-chip memory. The extra buffers and interfaces lead to an average increase of 2.19pps and 1.60pps in LookUp Tables (LUTs) and Flip-Flops (FFs) respectively, which is within the acceptable overhead. This shows the tradeoff between moving skip connections off chip and keeping them on chip, which is less efficient for small models i.e., with $c=2$. Table IV shows the throughput, IoU, energy and efficiency of the implemented models on the FPGA. All FPGA implementations achieved the same IoU as the corresponding software implementations.
V-C2 Edge Computing Platforms
@@
Figure 2: Pareto frontier showing the trade-off between accuracy (weighted IoU) and energy efficiency across different platforms and model sizes. FPGA on-chip: U-Net skip connections stored on chip, FPGA off-chip: U-Net skip connections stored off chip in the memory
@@ VI Conclusion and Future Work
In this work, we studied crack segmentation on resource-constrained edge platforms from two sides: The model design and the target hardware. For this, we scaled the U-Net architecture over three orders of magnitude in parameter count and analyzed the trade-off between IoU and computational cost. Our results show that reducing the model size can strongly lower the computational cost, while some of the smaller variants still remain close to the baseline in terms of segmentation accuracy. To improve our scaled U-Net models further, we applied knowledge distillation by using SegFormer B5 as the teacher model. This improved all U-Net variants.
We then evaluated all variants on several edge platforms and compared their throughput, power consumption, and energy efficiency. We also developed a custom FPGA-based architecture for single-batch inference. Among all tested platforms, the FPGA (on-chip) implementation with the U-Net Base 4 using KD and int8 quantization achieves an energy efficiency of 204.99 Frames/J and a throughput of 398 FPS, while improving the mean IoU by 2.79pps compared to the baseline model (U-Net Base 32 without KD). Compared to the Hybrid-Segmentor model from [12], which is - to the best of our knowledge - currently the best tested model on this dataset, our aforementioned U-Net Base 4 model with KD is 6.32pps higher in mean IoU by using the same train/val/test split.